ANNUAL REPORT OF THE CENTRAL BUREAU FOR ASTRONOMICAL TELEGRAMS

1996

During 1996 a total of 247 Circulars were issued, the most for any year since 1992. Not only is it encouraging to see the strong reversal of the downward trend in the recent annual totals, but it is good to be able to contradict the explanation previously given for supposing that such a trend would continue!

The main instrument for the change was the rather dramatic appearance of the brightest comet since 1976 and most spectacular comet to pass within 0.10 AU of the earth since 1556. C/1996 B2 (Hyakutake) was a fine precursor to 1997’s anticipated C/1995 O1 (Hale-Bopp), which is expected to have a more prominent dust tail and may well be brighter. Altogether, eleven discoveries of comets were reported in 1996, and there were five first-time recoveries of short-period comets. One of the discoveries had the orbit of an asteroid of the Themis family, but the presence of a persistent tail caused it to receive the cometary designation P/1996 N2 on Aug. 21. Five days earlier, there was the complementary announcement of 1996 PW, which with an orbit having a revolution period of 6000 years would normally suggest classification as a long-period comet—except that careful scrutiny showed it to be completely asteroidal in appearance!

Also contributing to the high activity in 1996 was the discovery of no fewer than 81 extragalactic supernovae, handsomely breaking the previous record from 1992. Two galactic novae were reported.

Other topics given particular coverage on Circulars were the source of hard x-rays GRO J1744−28 during the first half of the year and the γ-ray burst GRB 960720 during the second, each of them involving the discussion of optical candidates. There was also a suspected case of γ-ray bursts from the same source over a two-day interval.

The number of subscribers to the Computer Service and/or electronic version of the Circulars showed a healthy increase during the year, from 455 at the start to 567 at the end, but illicit redissemination continues, despite the appearance now of an explicit statement that the Circulars are copyright. As noted in last year’s report, that statement applies to the form of the publication itself, not to the information contained therein. Subscribers are welcome to pass on to their colleagues specific astronomical pieces of information gleaned from the Circulars, provided that they acknowledge the source, a point that is surely and simply a matter of courtesy. The number of subscribers to the printed version of the Circulars declined during the year from 443 to 419.

Most of the IAU Circulars during 1996 were prepared by Bureau Associate Director Daniel W. E. Green, the remainder by the undersigned. Responsibility for the administrative work of the Bureau has been shared by Donna Thompson and Muazzez Lohmiller. Gareth Williams has continued very effectively to maintain the Bureau’s presence on the World

Wide Web (URL http://cfa-www.harvard.edu/cfa/ps/cbat.html): this includes the free

availability of low-precision ephemerides for bright comets, Circulars more than a year old and even occasional current Circulars.

Brian G. Marsden Director of the Bureau